OpenStudy (anonymous):

The probability to win a prize is 0.25 with one go. The game is played for 10 times. X denotes the number of prices won. Find E(X)

OpenStudy (anonymous):

No calculator allowed.

OpenStudy (anonymous):

If I understand the problem correctly then, your x values will be {0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10}, and your p values will be?

OpenStudy (anonymous):

E(X) = np = 10(.25) = 2.5 This is a binomial distribution right?

OpenStudy (anonymous):

he plays the game 10 times - why the 0? or is 0, 1, 2... the number of prizes that he can possibly win?

OpenStudy (anonymous):

Yes, x represents the values of each outcome.

OpenStudy (anonymous):

oh and yea E(X) = np = mean = 2.5

OpenStudy (anonymous):

p(x) should be the probability of any particular outcome.

OpenStudy (anonymous):

we get to our second point. P(X<2) (X is less than or equal to two)

OpenStudy (anonymous):

expression for that probability is enough for answer no need to calculate it

OpenStudy (anonymous):

i liked at standardized variable list and it said 0.9773 for 2

OpenStudy (anonymous):

*looked at

OpenStudy (anonymous):

@Jire ...the second point: P ( X ≤ 2) ≈ 0.526 difficult to type full solution, not sure how to properly use equation box.

OpenStudy (anonymous):

you can just tell me how to basically get there

OpenStudy (anonymous):

It's just P(X=0) + P(X=1) + P(X=2)

OpenStudy (anonymous):

oh okay so using this formula gives me the - probability of that with 10 tries of 0.25 probability of getting the prize, less than or equal to two prizes are won??

OpenStudy (anonymous):

Well if it's P(X<2), then it's just P(X=0) + P(X=1)

OpenStudy (anonymous):

Yeah, you're interpreting it right.

OpenStudy (anonymous):

thx